show that the number of six letter words that can be formed out of letters of word assist in which s's alternate with each other is 12 .how?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear manisha

all  s can be placed at even places  or at odd places in = 2 ways

rest three places can be filled by =3*2*1 = 6 ways

so total ways = 2*6 =12

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin